OpenStudy (anonymous):

if sin^-1x+sin^-1y=2pi/3 then cos ^-1x+cos^-1y=

OpenStudy (anonymous):

1/sinx+1/siny=2pi/3 so sinx+siny=3/2pi

OpenStudy (anonymous):

sin(90-x)+sin(90-y)=3/2pi

OpenStudy (anonymous):

cosx+cosy=3/2pi

OpenStudy (anonymous):

1/cosx+1/cosy=2pi/3

ganeshie8 (ganeshie8):

First of all, \(\large \sin^{-1}x \color{Red}{\ne} \dfrac{1}{\sin x}\)

ganeshie8 (ganeshie8):

\(\large \sin^{-1}x\) is an inverse function, NOT a reciprocal, okay ?